What is the probability that a market inspector will discover, at least two, but fewer than four violations of the public health code?

1
Expert's answer
2022-02-08T12:05:19-0500

Use this table:


Solution:

"P(2\le x<4)=P(x=2)+P(x=3)=0.17+0.13=0.3"